OpenStudy (anonymous):

Factor Completely: 4x^2 − 81 (4x − 9)(x + 9) (2x + 9)(2x + 9) (2x + 9)(2x − 9) (2x − 9)(2x − 9)

OpenStudy (ankit042):

(a^2-b^2) = (a+b)(a-b) Now you can apply this and get your answer!

OpenStudy (uri):

Y U NO TRY TO DO IT..OMGG

Directrix (directrix):

This problem illustrates one of the basic factoring patterns: the difference of two squares. 4x^2 − 81 = (2x)² - (9)² = ?